A ten-year bond has a yield of 13% and a duration of 7.209 years. If the bond's yield increases by 50 basis points, what is the percentage change in the bond's price as predicted by the duration formula? (Input the value as a positive value. Do not round intermediate calculations. Round your answer to 2 decimal places.) The bond's price
Answer #1

Case 1: Semiannual coupon payments
% change=-7.209/(1+13%/2)*0.50%=-3.38%

Case 2: Annual coupon payments
% change=-7.209/(1+13%)*0.50%=-3.19%